### v1

# WaveGuard: Robust Deepfake Detection and Source Tracing via Dual-Tree Complex Wavelet and Graph Neural Networks

Abstract
Deepfake technology has great potential in the field of media and entertainment, but it also brings serious risks, including privacy disclosure and identity fraud. To counter these threats, proactive watermarks methods have become a research hotspot by embedding invisible signals to build active protection schemes. However, existing methods are vulnerable to watermark destruction under malicious distortions, which leads to insufficient robustness. Moreover, embedding strong signals may degrade image quality, making it challenging to balance robustness and imperceptibility. To address these issues, this paper proposes a proactive watermarking framework called WaveGuard, which explores frequency domain embedding and graph-based structural consistency optimization. In this framework, the watermark is embedded into the high-frequency sub-bands by dual-tree complex wavelet transform (DT-CWT) to enhance the robustness against distortions and deepfake forgeries. By leveraging joint sub-band correlations and utilizing specific sub-band combinations, the framework performs robust extraction for source tracing and semi-robust extraction for deepfake detection. In addition, the framework adopts dense connectivity strategies to promote feature reuse and improve robustness under complex attacks. To enhance imperceptibility, we propose a Structural Consistency Graph Neural Network (SC-GNN), which constructs graph representations to extract structural features from the original and watermarked images, ensuring consistency between them and reducing perceptual artifacts to improve visual quality. Finally, an attention mechanism named Tanh-based Spatial Embedding Attention Module (TSEAM) is designed to further refine global and local features, thereby improving watermark concealment without compromising robustness. Experimental results show that the proposed method performs exceptionally well in face swap and face replay tasks. The code will be published at https://github.com/vpsg-research/WaveGuard.
Index Terms:
Deepfake Detection, Source Tracing, Frequency-domain Embedding, Graph Neural Network (GNN).
I Introduction
Fig. 1: Three categories of approaches for balancing robustness and invisibility in proactive deepfake detection.
The rapid development of artificial intelligence generated content (AIGC) has greatly improved the ability of automation and innovation. However, its wide application also brings risks, especially the deepfake technology, which can generate forged media, threaten privacy, encourage identity theft and spread false information[1]. Deepfake generation techniques mainly includes face swapping and face reenactment. Face swapping[2] replaces the identity features of the target into the source image while preserving the background and semantic information. On the other hand, face reenactment[3, 4] manipulates expressions or poses to map dynamic features onto the target image while maintaining identity consistency. Although deepfake has positive application in some fields, its abuse has caused serious social concern. Thus, it is very important to develop effective preventive measures.
Currently, most deepfake detection methods rely on passive detection, which determines the authenticity of images by analyzing anomalous features present in forged content. Typical methods include spatial domain artifact detection, such as color inconsistency, abnormal boundary mixing and irregular facial muscle movement[5, 6]. Other approaches explore frequency-domain features, identifying forgery clues like spectral distortions and encoding residuals[7]. In addition, some work also introduces biometric consistency (such as blink frequency and heartbeat signal) as auxiliary information to enhance the identification ability of forged content[8]. Although these method are effective in content authenticity verification, it usually comes into play after the generation and dissemination of forged content, and lacks effective traceability and interpretability.
Recently, the concept of proactive approach has emerged as a preventive strategy, and its core idea is to embed invisible signals when the original content is released. This approach is mainly divided into proactive distortions[9, 10, 11, 12, 13] and proactive watermarks[14, 15, 16, 17, 18, 19, 20, 21, 22]. The former interferes with the generation process of a deepfake model by embedding invisible adversarial perturbation in the original image, destroying or counteracting the effect of the deepfake model, and making the generated content appear visually abnormal. While this provides early protection, it lacks traceability and verification capabilities. In contrast, the proactive watermarks method is to embed the watermark information in the original content in advance, and then realize the dual functions of traceability and detectability by detecting the absence or change of the watermark. This method ensures copyright protection and source tracing, especially avoids the unexplained problems in the passive deepfake detection results. However, the existing proactive watermarks methods face challenges in robustness, particularly when dealing with complex deepfake attacks. As deepfake technology advances, the forgery techniques have become more subtle and sophisticated, making it increasingly difficult to extract watermarks effectively. Meanwhile, maintaining invisibility is equally important. Improving robustness often requires stronger embedding, which may introduce perceptual artifacts. Conversely, enhancing invisibility will reduce the robustness against distortion. Thus, balancing robustness and invisibility has become a key challenge for proactive watermarks.
To better understand existing approaches for balancing robustness and invisibility, Fig. 1 presents three different categories of deep learning-based watermarking frameworks. The first two cases adopt spatial domain embedding strategy, and hide watermark information by directly modifying image pixel values. Although the spatial domain method is simple and efficient, it still faces significant challenges when confronted with complex attacks. Because pixel modification inevitably damages the embedded watermark, this method makes the spatial watermark highly sensitive to distortion and deepfake attack. Although Case1 and 2 have shortcomings in robustness, they respectively represent two typical strategies to improve the invisibility of watermark. In Case1, Generative Adversarial Network (GAN)[23] generate high-quality watermarked images through adversarial training, which makes the watermark image visually similar to the original image to enhance invisibility. However, GANs often face problems of unstable training and mode collapse, which result in inconsistent and unstable outputs. In Case2, Variational Autoencoder (VAE)[24] extract latent features from both the original and watermarked images and optimize their similarity through a loss function, aiming to enhance the watermarked image quality while preserving natural image structures. However, due to the constraints of latent space, the reconstructed images generated by VAEs are often blurred and the detailed information is easily lost. Especially when different datasets are pretrained, the effect of feature extraction is not satisfactory.
To address the limitations of existing methods in both robustness and invisibility, we propose a dual-channel watermarking framework called WaveGuard. It integrates frequency-domain embedding with a graph neural network-based structural consistency constraint to support robust source tracing and deepfake detection. Aiming at the problem that the local details are often lost and the directional texture disturbance often leads to the invalidation of the watermark, this framework innovatively uses the translation invariance and directional selectivity of the dual-tree complex wavelet transform (DT-CWT)[25] to embed the watermark into the high-frequency sub-band, thus significantly enhancing its robustness to common distortions and Deepfake forgery. In addition, the traceability and detection function of the watermark are realized by using the correlation between the joint sub-bands. By using different sub-band combinations, ensuring robustness for traceability and semi-robustness for detection. To further enhance performance, the framework adopts dense connection strategies to promote efficient feature reuse and improve the robustness of the watermark. Additionally, in order to solve the problem that traditional watermarking methods introduce obvious artifacts in the visual quality of images, this paper designs a structural consistency graph neural network (SC-GNN), as shown in Case 3 in Fig. 1. Unlike VAE, which extracts the latent features of images, SC-GNN captures the spatial correlation of original images and watermarked images by constructing graphical representations. The graph convolution network (GCN) further extracts and optimizes the structural features, which ensures that the watermark image maintains the integrity of the original image and reduces the perceptual differences. To further enhance imperceptibility, the framework employs a Tanh-based Spatial Embedding Attention Module (TSEAM) to preserve global structures and refine local details, thereby improving the visual quality of watermarked images. Our contributions can be summarized as follows:
• We adopt a frequency-domain watermarking strategy using DT-CWT, leveraging high-frequency sub-bands to enhance resistance against distortions and forgeries. By utilizing joint sub-band correlations for robust and semi-robust embedding, the method ensures reliable source tracing and deepfake detection.
• We propose a Structural Consistency Graph Neural Network (SC-GNN), which leverages the internal structural relationships of the image to enhance the imperceptibility of the watermarked image. This method constructs graph representations of both the original and watermarked images, capturing spatial correlations to better maintain visual consistency.
• We construct an end-to-end dense connection network WaveGuard, which improves the flexibility and resilience of watermark embedding by optimizing feature transmission and reuse. Experimental results validate the advantages of our method in watermark recovery and deepfake detection, achieving superior robustness and imperceptibility compared to existing approaches.
Fig. 2: Two-level DT-CWT decomposition.
II RELATED WORKS
II-A Proactive Distortions
Proactive defense strategy destroys or counteracts the generation effect of a deepfake model by embedding invisible adversarial perturbation in the original image. This method provides preventive protection before the image is released, ensuring that when the image is maliciously used for deepfake, the generated content will appear visually abnormal. For example, Ruiz et al.[9] proposed a confrontational attack method that destroyed the conditional image translation network through invisible perturbations to prevent deepfake face manipulation. Similarly, CMUA[10] significantly reduces the forgery ability of several deepfake models by generating antagonistic perturbations. In addition, TAFIM[11] uses an attention mechanism to fuse specific perturbations and generate perturbations that cannot be identified by the model, thereby preventing image manipulation. Furthermore, Anti-Forgery[12] interferes with the generation of deepfake content by introducing perceptual perturbations into the Lab color space, creating significant artifacts and maintaining robustness against input transformations. At the same time, DF-RAP[13] provides robust image protection by modeling the social network compression mechanism to defend against deepfake. Although proactive defense can interfere with the generation of deepfake at an early stage, thus providing some protection, it still has some key limitations. For example, powerful generation models may bypass or neutralize these perturbations, thus weakening their effectiveness. In addition, proactive distortions lacks tracking ability and post-verification, and it is difficult to judge whether the image has been tampered with after publication.
II-B Proactive Watermarks
Digital watermarking is a technology to protect copyrights, verify integrity, or source tracing by embedding hidden information in digital media. HiDDeN[26] is the first end-to-end trainable digital robust watermarking framework. It uses the encoder-decoder structure of deep neural networks to hide information and achieves efficient recovery under various image distortions, opening a new research direction in the field of digital watermarking. Aiming at the problem of non-differentiable JPEG compression, MBRS[27] further enhances the robustness of the model by randomly switching real JPEG, simulated JPEG, and noise-free layers in small batches. Similarly, CIN[28] integrates a reversible network and an attention module to enhance robustness while maintaining high invisibility and resistance to distortions. To overcome the limitations of traditional single-stage end-to-end frameworks, TSDL[29] employs a two-stage design to strengthen watermark robustness and improve model stability and convergence speed. Further, ARWGAN[30] enhances both robustness and invisibility through feature fusion and attention mechanisms, aiming to achieve low bit error rates while preserving high visual quality. In addition, Ma et al.[31] constructed a robust watermark model using Swin Transformer and Deformable Convolutional Network (DCN), and proposed a pattern transfer noise layer to improve generalization performance in multi-distortion scenarios. For cross-media distortion (such as print scanning, print camera and screen camera, etc.), researchers have proposed various solutions, including LFM[32], Stegstamp [33], RIHOOP [34], PIMoG [35], GDS-WM [36], and WRAP[37]. These methods target different distortion types by designing optimized strategies to ensure stable watermark extraction and verification after physical printing and re-digitization.
In the emerging field of proactive face forensics, existing methods are limited. Yu et al. [14, 15] proposed an artificial fingerprint embedded within generative models for deepfake detection and source tracing, demonstrating strong scalability and robustness against complex noise. FaceSigns[16] differs from traditional robust watermarking by emphasizing robustness under common distortions while selectively failing under malicious distortions, thus achieving more accurate deepfake detection. Similarly, FaceGuard[17] achieves proactive deepfake detection by embedding semi-robust watermarks. Building on these efforts, Sepmark[18] proposed a separate watermarking framework that effectively addresses both common and malicious distortions by combining robust and semi-robust watermarking using a single encoder and separable decoder. At the same time, Wu et al.[19] proposed a new framework that adjusts robust watermarks into adversarial ones to improve detection and tracing capabilities. It also supports passive detection, making it easier to combine proactive and passive detection. Wang et al. [20] proposed a robust watermarking framework based on identity semantics, achieving deepfake detection and traceability with strong robustness and accuracy. In addition, Lampmark[21] uses an untrained pre-defined pattern watermark to actively defend against deepfake, combining structural sensitivity and robustness to achieve efficient detection. With the growing use of multifunctional watermarking, Dual Defense[22] introduced an invisible, robust watermarking method that combines interference and traceability to prevent effective face-swapping outputs and allow copyright tracking.
Overall, although various methods have explored watermark robustness and deepfake detection, there are still challenges in improving adaptability and imperceptibility. Some methods rely on spatial domain embedding and are easily affected by tampering, while others lack effective visual consistency maintenance mechanism. In contrast, our method integrates frequency domain embedding and structural consistency optimization, and achieves the unity of forgery detection and source tracking while taking into account robustness and invisibility.
III METHODOLOGY
III-A Preliminaries
III-A1 Dual-tree complex wavelet transform
Dual-Tree Complex Wavelet Transform (DT-CWT) is a wavelet transform with translation invariance and limited redundancy, with a redundancy factor of 2. Unlike traditional DWT, DT-CWT decomposes the image into six directional high-frequency sub-bands, corresponding to angles of $15^{\circ}$, $45^{\circ}$, $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$. This allows it to accurately capture the directional features and details of the image. As shown in Fig. 2, DT-CWT adopts a circular layout to illustrate the directional and hierarchical structure of two-level decomposition. The first-level decomposition splits the image into six high-frequency sub-bands ($U^{H}_{1,d}$, where $d$ takes values from 1 to 6) and one low-frequency sub-band ($U^{L}$ at the center). The second-level decomposition further splits the first-level low-frequency sub-band, generating six new high-frequency sub-bands ($U^{H}_{2,d}$) and one low-frequency sub-band, with the size of each sub-band halved. Blue arrows represent the first-level high-frequency sub-bands, and red arrows represent the second-level high-frequency sub-bands, with consistent directions across both levels. DT-CWT possesses multi-scale and multi-directional analysis capabilities, which precisely align with the need for accurate characterization of texture boundaries and directional details in watermarking. In the face of compression, blurring, and sophisticated deepfake forgeries, its approximate translation invariance offers a stable and reliable embedding space in the frequency domain, thereby enhancing the robustness of watermark signals against fine-grained distortions.
Fig. 3: Frequency response of four sub-bands with watermark embedding and propagation. Sub-band 1 and sub-band 3 (pink) represent the watermark embedding regions, while sub-band 4 and sub-band 6 (blue) indicate the propagation of watermark information due to sub-band redundancy.
III-A2 Joint Sub-bands
Joint sub-band correlation was first proposed by Huan et al.[38], who significantly enhanced the robustness of video watermarking through the directional selectivity and translation invariance of DT-CWT. To further verify its role in image watermark embedding, we designed a simple experiment. A standard grayscale image (256 × 256 pixels) was decomposed by DT-CWT to obtain a low-frequency sub-band and six high-frequency sub-bands. Watermarks were embedded in sub-band 1 and sub-band 3. The results show that, although the watermark is only embedded in sub-band 1 and sub-band 3, sub-band 4 and sub-band 6 carry some watermark information due to the redundancy between sub-bands. The correlation with the original watermark is significantly high. Fig. 3 illustrates the distribution of the watermark in the frequency domain: pink represents the watermark embedding area, blue indicates the propagation area, and arrows show the propagation path. There is natural redundancy between the joint sub-bands, which enables the watermark information to spread in multiple high-frequency sub-bands, so it still has strong recovery ability under local tampering and frequency disturbance. Based on the correlation design of joint sub-bands, watermark information can spread among multiple sub-bands, thus maintaining stable extraction in different attack scenarios, which has become the key supporting mechanism for the system to achieve robust traceability and semi-robust deepfake detection.
III-B Model Architecture
In order to realize a robust watermark embedding and detection system for deepfake, this paper designs a framework, which mainly consists of the following five parts: (1) Frequency Subband Selection Module (FSSM): The frequency subband selection module converts the input image from the spatial domain to the frequency domain and extracts specific frequency sub-band information. (2) Encoder (En): The encoder uses a deep learning network to fuse the specific sub-band information output by the frequency subband selection module with the watermark information, generating the embedded image. (3) Noise Pool Layer (NPL): Simulates common distortions and malicious distortions in actual scenarios to generate the distorted image. (4) Separable Decoder: Includes DensTracer and DensDetector, which are used to extract watermarks under distorted conditions to realize image traceability and detect deepfake, respectively. (5) Structural Consistency Graph Neural Network (SC-GNN): By comparing the image characteristics before and after watermark embedding, the watermark embedding process is optimized to ensure the visual consistency of the watermark image and improve the invisibility of the watermark.
Fig. 4: Overall Architecture of WaveGuard. Frequency Subband Selection Module ($FSSM$) extracts and processes high-frequency sub-bands in the frequency domain, providing essential features for watermark embedding, tracing, and detection. The encoder ($En$) embeds the watermark into the high-frequency sub-bands and reconstructs the watermarked image $I_{w}$. The Noise Pool Layer ($NPL$) applies various distortions to $I_{w}$ to generate distorted images $I_{n}$. The tracer decoder ($Tr$) extracts the watermark $w_{t}$ from $I_{n}$ to evaluate its robustness, while the detector decoder ($De$) extracts the correct watermark $w_{d}$ under common distortions and outputs random-like messages under malicious distortions. In addition, Structural Consistency Graph Neural Network ($SC-GNN$) extracts the structural features of the original image $I_{o}$ and the watermarked image $I_{w}$ by constructing graph representation to ensure the consistency between them.
III-B1 Frequency Subband Selection Module (FSSM)
The input image is first decomposed into three channels: Y, U, and V, with the U channel as the main processing object. Secondly, the U channel is decomposed in the frequency domain using the dual-tree complex wavelet transform (DT-CWT), obtaining a low-frequency sub-band (low_pass) and high-frequency sub-bands (high_pass). High-frequency sub-bands contain frequency information in six directions. By selecting sub-bands 1 and 3, key high-frequency details are extracted for watermark embedding, while sub-bands 4 and 6 are reserved for the subsequent traceability stage. This module completes the frequency domain conversion and sub-band selection, providing input for watermark embedding and extraction to the encoder and two separable decoders.
III-B2 Encoder (En)
The high-frequency sub-band feature output extracted by the frequency subband selection modul undergo Conv-BN-ReLU processing in the Spatial Feature Extraction Module, generating a refined feature map that captures both spatial and fine-grained details. The watermark is transformed into a two-dimensional matrix via a linear layer, resized to match the feature map dimensions, and processed by a convolutional block to extract watermark-specific features. These features are then fused with the shallow spatial features. A three-stage dense connection structure is used for hierarchical watermark embedding. At each stage, the feature map and watermark are concatenated and processed through the DenseBlock Module, which consists of a 1×1 convolution for dimensionality reduction and a 3×3 convolution for feature enhancement. In this process, two different attention mechanisms are adopted: Tanh-based Spatial Embedding Attention Module (TSEAM) and Spatial-wise Image Attention Module (SIMAM). TSEAM is improved on the basis of Convolutional Block Attention Module, which acts on the whole network structure to highlight the important areas in the image. By replacing the ReLU activation function with tanh, the continuity and integrity of watermark features in the [0,1] interval are better preserved. SIMAM is applied after each DenseBlock to weight the local feature map and help the model capture the image details, thus improving the invisibility of the watermark and the robustness to interference. The combination of TSEAM and SIMAM optimizes the watermark embedding process from two dimensions of global features and local details, which not only ensures the effective use of global semantic information, but also realizes the precise control of local features, thus improving the overall visual quality and embedding stability. After the three-stage fusion, the encoder compresses the features using a Down Block, reducing their spatial dimensions while preserving critical information. The final compressed features generate the watermark difference map (wm_diff), which is added to the input image to complete the watermark embedding process. watermark embedding is achieved using the following formula:
$\boldsymbol{U_{\mathit{embed}}=U+\alpha\cdot\mathit{attention\_weight}\cdot\mathit{wm}_{\mathit{diff}}}$ (1)
In Formula (1), $U$ represents the original high-frequency sub-band, $\alpha$ controls the watermark embedding strength, attention_weight is the mask generated by the attention mechanism to weight the watermark embedding area, and $\text{wm}_{\text{diff}}$ is the generated watermark difference map. Through this process, the watermark information is embedded into the original high-frequency sub-band, forming a watermarked high-frequency sub-band $U_{\text{embed}}$. The post-processing module uses the inverse dual-tree complex wavelet transform (IDTCWT) to reconstruct the watermarked U-channel, merges it with the Y and V channels to form watermarked YUV images, and converts them to RGB images $I_{w}$, completing the embedding process.
III-B3 Noise Pool Layer (NPL)
In practical applications, when images are uploaded or spread through social media, the quality is often degraded due to common processing operations such as compression and blurring or noise interference. To this end, we have established two operation sets: Common Pool and Malicious Pool, which are used to assist framework training. Among them, Common Pool contains common benign distortions in daily use, while Malicious Pool introduces deepfake model to improve the robustness of watermark in response to malicious attacks.
III-B4 Separable Decoder
The two decoders adopt a DenseNet structure similar to that of the encoder. The decoder takes the high-frequency sub-band embedded with the watermark as input, generates the initial feature map through the feature extraction module, and gradually fuses the decoded information through the multi-layer DenseBlock module. Finally, the high-dimensional features are compressed by the down-sampling module to generate a single-channel watermark difference map, which is adjusted to the target size through convolution and interpolation, and finally restored to one-dimensional watermark information.
Tracer and Detector share the same structure in design but differ in sub-band selection and target tasks. To enhance robustness, Tracer uses the 1st, 3rd, 4th, and 6th high-frequency sub-bands to extract features and handle various distortions, achieving watermark traceability with the goal of $W_{t}=W_{o}$. The Detector decoder only uses the 1st and 3rd sub-bands to extract features for the semi-robust requirement of forgery and tampering detection. Its target is $W_{d}=W_{o}$ for normal distortions and $W_{d}\neq W_{o}$ for malicious distortions. This difference in sub-band selection ensures that the two decoders meet the requirements for robust extraction and efficient detection, respectively.
III-B5 Structural Consistency Graph Neural Network (SC-GNN))
First, the original image and the watermarked image will be sent to the Graph Construction Module. In this module, the pixels of each image are regarded as nodes in the graph, and the edges of the graph are constructed by 4- neighborhood connection (upper, lower, left and right adjacent pixels), thus generating graph representations of the original image and the watermarked image. In the graph construction module, each pixel of the image corresponds to a node, and the connection relationship (edge) between nodes is defined by the spatial relationship between adjacent pixels. Then, the graph data is sent to the graph convolution network (GCN). By capturing the spatial dependence between nodes, GCN extracts global and local features, optimizes the structural consistency between the watermark image and the original image, and reduces perceptual artifacts, thus improving the invisibility and robustness of the watermark. In the training process, the structural consistency loss function is used to further ensure the structural consistency between the watermark and the original image.
III-C Loss Functions
To optimize our frequency-domain watermarking framework for embedding, tracing, and detection, we define loss functions to optimize the encoder-decoder framework. $\theta$ represents trainable parameters, with $\theta_{co}$ and $\theta_{ma}$ denoting common and malicious distortions sampled from the Noise Pool Layer (NPL). The details are as follows:
1. Embedding Loss
The goal of the embedding loss is to ensure that the watermarked high-frequency sub-band $U_{\text{embedded}}$ is highly similar to the original $U$-channel. It is defined as:
$\mathcal{L}_{En}=\|U-E_{n}(FSSM(I_{o}),W_{o})\|_{2}^{2}=\|U-U_{embedded}\|_{2}^{2}$ (2)
where $FSSM(I_{o})$ represents the original image processed by the frequency subband selection module, and $W_{o}$ is the embedded watermark.
2. Tracing Loss
The tracing loss ensures that the robust decoder $Tr$ can accurately extract the embedded watermark $W_{t}$ from the watermarked image $I_{w}$, even after various distortions. It is defined as:
$\mathcal{L}_{Tr}=\|W_{o}-Tr(\theta,NPL(\theta_{co}+\theta_{ma},I_{w}))\|_{2}^{2}=\|W_{o}-W_{t}\|_{2}^{2}$
where $NPL$ is the noise processing layer that simulates distortions.
3. Detection Loss
The detection loss ensures that the semi-robust decoder $De$ behaves differently when facing common and malicious distortions:
The extracted watermark $W_{d}$ from the commonly distorted image $I_{no}$ should match the original watermark $W_{o}$:
$\mathcal{L}_{De1}=\|W_{o}-De(\theta,NPL(\theta_{co},I_{w}))\|_{2}^{2}=\|W_{o}-W_{d}\|_{2}^{2}$
For malicious distortions, the output $W_{d}$ of the decoder $De$ should approach random guessing (represented as 0):
$\mathcal{L}_{De2}=\|0-De(\theta,NPL(\theta_{ma},I_{w}))\|_{2}^{2}=\|0-W_{d}\|_{2}^{2}$
4. SC-GNN Loss
The SC-GNN loss optimizes the structural consistency between the original image and the watermarked image. It is defined as:
$\mathcal{L}_{\text{GNN}}=\bigl{\|}\text{GCN}(\text{GCM}(I_{o}))-\text{GCN}(\text{GCM}(I_{w}))\bigr{\|}_{2}^{2}=\|f_{o}-f_{w}\|_{2}^{2}$ (6)
where $\text{GCM}(I_{o})$ and $\text{GCM}(I_{w})$ denote the graph representations constructed from the original and watermarked images, respectively, using the Graph Construction Module (GCM). These graphs are subsequently processed by the Graph Convolutional Network (GCN) to extract structural features, denoted as $f_{o}=\text{GCN}(\text{GCM}(I_{o}))$ and $f_{w}=\text{GCN}(\text{GCM}(I_{w}))$. The SC-GNN loss minimizes the mean squared error (MSE) between these feature representations to enforce structural consistency.
5. Total Loss
The total loss is defined as a weighted combination of the above loss components, balancing their contributions during training. It is expressed as:
$\mathcal{L}_{Total}=\lambda_{1}\mathcal{L}_{GNN}+\lambda_{2}\mathcal{L}_{En}+\lambda_{3}\mathcal{L}_{Tr}+\lambda_{4}\mathcal{L}_{De1}+\lambda_{5}\mathcal{L}_{De2}$ (7)
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}$ are hyperparameters that control the weights of each loss component.
Fig. 5: Illustrates the visual effects of operations performed on watermarked images. The original image $I_{o}$ and the watermarked image $I_{w}$ are shown in the first and second rows, respectively. The results of applying different operations to the watermarked image, denoted as $I_{n}$, are displayed in the third row. The three columns on the right showcase the effects of deepfake operations, while the remaining columns present the results of benign operations. The fourth row shows the watermarked residual signals, represented as $|I_{o}-I_{w}|$, which are processed to enhance the visualization of differences. All images are sized $128\times 128$.
IV EXPERIMENTS
IV-A Implementation Details
Datasets. This research is primarily based on the high-quality CelebA-HQ[39, 40]dataset, which contains 30,000 facial images and 6,217 unique identities. According to the official partitioning, it is divided into training, validation, and testing sets, where the training set contains 24,183 images, and the validation and test sets contain 2,993 and 2,824 images, respectively. To verify the adaptability and robustness of the algorithm across different resolutions and simulate the diverse requirements of practical application scenarios, the experiment uses two types of images with resolutions of 128×128 and 256×256 for testing.
Manipulation Pools. In order to comprehensively evaluate the robustness and adaptability of the algorithm under different distortion conditions, this study constructs two types of operation pools: the Common Pool and Malicious Pool. The Common Pool includes a variety of common image distortion operations, such as {Identity, JpegTest, Resize, GaussianBlur, MedianBlur, Dropout, SaltPepper, GaussianNoise} . These operations are used in the training and testing stages to simulate the general distortions that may occur in real-world scenarios, while ensuring that the facial content of the image remains unchanged. The Malicious Pool includes two common types of deepfake technologies: face swapping and face reenactment. Specifically, SimSwap[2] is used for face swapping, while GANimation[4] and StarGAN[3] are used for face reenactment. In the SimSwap task, the target face is randomly selected from the CelebA verification set. ”In the tasks of GANimation and StarGAN, the target expressions and attributes are randomly selected. The attributes are chosen from the set Male, Young, BlackHair, BlondHair, BrownHair. All forged images are generated using the publicly released pre-trained models to ensure the high credibility and diversity of deepfake images.
Table I
OBJECTIVE VISUAL QUALITY EVALUATION OF WATERMARKED IMAGES ON CELEBA-HQ: MODEL, IMAGE SIZE, MESSAGE LENGTH, PSNR, AND SSIM
Model Image Size Message Length PSNR ↑ SSIM ↑
MBRS[27] $128\times 128$ 30 33.0456 0.8106
CIN[28] $128\times 128$ 30 42.4135 0.96286
PiMoG[35] $128\times 128$ 30 37.7271 0.9470
SepMark[18] $128\times 128$ 30 38.5112 0.9588
Ours $128\times 128$ 30 47.5555 0.9957
FaceSigns[16] $256\times 256$ 128 32.3319 0.9211
SepMark[18] $256\times 256$ 128 38.5646 0.9328
Ours $256\times 256$ 128 43.3082 0.9893
Parameters. This research is based on PyTorch[41] and runs on an NVIDIA RTX 4090 GPU. The Adam optimizer[42]is initially set with a learning rate of 0.0005, which is later dynamically adjusted during training. The total loss function is composed of multiple loss terms, which are combined by weighting to balance the influence of each loss on model optimization. As shown in Formula (7), during the training process, the weight parameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to 0.01, 1, 10, 10, and 10, respectively. These weights are adjusted reasonably to reflect the relative importance of each loss term and ensure the convergence and stability of the model in the optimization process.
Table II
COMPARISON OF BIT ERROR RATE IN WATERMARK RECOVERY FOR CELEBA-HQ ACROSS COMMON AND DEEPFAKE DISTORTIONS AT 128$\times$128 AND 256$\times$256 RESOLUTIONS
$128\times 128$ $256\times 256$
Distortion MBRS[27] CIN[28] PIMoG[35] SepMark[18] Ours FaceSigns[17] SepMark[18] Ours
Tracer Detector Tracer Detector Tracer Detector Tracer Detector
Identity 0.0000% 0.0000% 0.0366% 0.0000% 0.0000% 0.0000% 0.0000% 0.0136% 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.2597% 2.7514% 19.5562% 0.2136% 0.2172% 0.0000% 0.0000% 0.8258% 0.0075% 0.0069% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0767% 0.0059% 0.0212% 0.0000% 0.0000% 1.0726% 0.0000% 0.0000% 0.0000% 0.0000%
GaussianBlur 0.0000% 22.7786% 0.1169% 0.0024% 0.0035% 0.0018% 0.0021% 0.1671% 0.0000% 0.0274% 0.0000% 0.0056%
MedianBlur 0.0000% 0.0307% 0.0992% 0.0012% 0.0012% 0.0013% 0.0017% 0.0977% 0.0000% 0.0000% 0.0000% 0.0000%
Dropout 0.0000% 0.0000% 0.4828% 0.0000% 0.0000% 0.0000% 0.0000% 17.5615% 0.0058% 0.0000% 0.0018% 0.0027%
SaltPepper 0.0000% 0.0378% 2.3667% 0.0413% 0.0106% 0.0132% 0.0156% 12.3238% 0.0008% 0.0003% 0.0077% 0.0045%
GaussianNoise 0.0000% 0.0000% 12.7396% 0.7460% 0.8735% 0.0211% 0.0327% 7.0697% 0.0578% 0.0622% 0.0039% 0.0045%
Average 0.0325% 3.1998% 4.4408% 0.1329% 0.1409% 0.0047% 0.0065% 12.8914% 0.0249% 0.0121% 0.0017% 0.0023%
SimSwap 19.3744% 48.5068% 8.6745% 13.8255% 50.8829% 0.0287% 46.7845% 49.9463% 7.9068% 45.9117% 0.0086% 44.5411%
GANimation 0.0000% 0.0000% 0.4802% 0.0000% 36.7938% 0.0000% 46.7879% 45.4172% 0.0020% 43.8524% 0.0000% 44.5431%
StarGAN (Male) 18.3133% 67.2568% 9.2044% 0.1157% 52.6003% 0.0051% 46.7828% 50.2617% 0.0033% 45.4624% 0.0011% 44.5395%
StarGAN (Young) 17.0562% 69.0805% 8.7465% 0.1074% 52.3678% 0.0019% 46.7828% 50.3649% 0.0030% 45.5319% 0.0008% 44.5458%
StarGAN (BlackHair) 19.2233% 58.7913% 10.5312% 0.1416% 49.2434% 0.0022% 46.7845% 50.2576% 0.0019% 45.1299% 0.0022% 44.5403%
StarGAN (BlondHair) 18.4478% 72.9733% 10.3506% 0.1712% 48.6084% 0.0013% 46.7811% 50.9829% 0.0022% 44.6953% 0.0009% 44.5415%
StarGAN (BrownHair) 17.6381% 79.5857% 9.0675% 0.0980% 52.3820% 0.0064% 46.7828% 50.7884% 0.0017% 45.9950% 0.0014% 44.5414%
Average 15.7219% 56.5992% 8.1507% 2.0656% 48.9827% 0.0065% 46.7838% 49.7170% 1.1316% 45.2255% 0.0021% 44.5418%
Baselines. In order to verify the performance of this research method, we selected several representative watermarking and proactive watermarks frameworks as baselines. Although this research belongs to the field of proactive watermarks, due to the limited existing methods and the scarcity of open source codes, we also refer to the robust watermarking method to evaluate its stability and extraction performance. In the watermarking domain, three robust watermarking methods were selected: MBRS [27], CIN [28], and PIMOG [35] (suitable for 128×128 images). In the proactive watermarks domain, we include FaceSigns[16] (for 256×256 images) and SepMark[18] (which supports both 128×128 and 256×256 resolutions). All the other baseline methods use publicly available pre-trained model weights to ensure fairness.
IV-B Visual Quality
To evaluate the visual quality of the image after watermark embedding, this paper adopts two indicators: average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM)[43], which respectively quantify the changes in image signal distortion and structural fidelity, fully reflecting the impact of embedding on the image. Table 1 quantitatively verifies that our method has the best visual quality in all baselines and shows better invisibility. Experimental results show that our method achieves the highest PSNR and SSIM values at resolutions of 128×128 and 256×256. This improvement is primarily attributed to SC-GNN and attention-enhanced embedding. In addition, Figure 5 verifies that there is almost no difference between the watermark image and the original image from a subjective point of view. Combining the quantitative evaluation in Table 1 with the visual comparison in Fig. 5, we can see that our watermark has superior invisibility in both objective indicators and subjective perception.
IV-C Robustness Test
The average bit error rate (BER) is selected as the main evaluation index in the robustness test. In any case, the BER value of Tracer should be close to 0%, and the BER value of the semi-robust Detector should be close to 50% when it encounters malicious distortion. In the case of common distortion, the BER value should be close to 0%.
$\text{BER}(W_{\text{rec}},W_{o})=\frac{\sum_{i=0}^{l-1}\left|W_{\text{rec}}^{i}-W_{o}^{i}\right|}{l}$ (8)
where $W_{o}$ is the original watermark, $W_{\text{rec}}$ represents the recovered watermark, and $l$ denotes the length of the binary watermark sequence. The BER calculates the ratio of error bits to the total length of the watermark, serving as a direct measure of recovery accuracy under various distortion scenarios.
Table 2 quantitatively compares the BER (Bit Error Rate) performance of different methods under common distortion and deepfake forgery, covering 128×128 and 256×256 resolutions. Experimental results show that our method achieves the lowest BER in all test scenarios, showing excellent robustness. Under ordinary distortion, the BER of Tracer and Detector are both close to 0%, which ensures that the watermark can be recovered with high precision under all kinds of interference. However, the BER of the contrast method is obviously higher under various common distortion conditions, which indicates that its robustness is weak. Under the deepfake forgery, the BER of Tracer is still close to 0%, while the BER of semi-robust Detector is maintained close to 50%, which effectively distinguishes the true and forged contents. However, the existing methods have large fluctuations in BER and lack of detection reliability. This robustness is mainly due to the frequency domain embedding strategy. Compared with the spatial domain method, the frequency domain watermark is more stable under common distortion, ensuring the integrity of the watermark information. At the same time, the design of joint sub-band enhances the discrimination ability of semi-robust Detector under deepfake forgery, so that it can stably distinguish real and forged content.
Table III
ABLATION RESULTS FOR PSNR, SSIM, AND TRACER’S BIT ERROR RATE (%), WHERE ‘×’ DENOTES NON-USAGE AND ‘$\sqrt{}$’ DENOTES USAGE.
Case1 Case2 Case3 Case4 Proposed
Joint Sub-bands $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Attention module $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
frequency domain $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
SC-GNN $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
PSNR $\uparrow$ 45.2321 43.7024 49.0436 42.8801 47.5555
SSIM $\uparrow$ 0.9956 0.9823 0.9955 0.9517 0.9957
Identity 0.0000% 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.0000% 0.0000% 0.0198% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0057% 0.0000% 0.0000%
GaussianBlur 0.0202% 0.0021% 0.0029% 0.0021% 0.0018%
MedianBlur 0.0021% 0.0017% 0.0011% 0.0005% 0.0013%
Dropout 0.0008% 0.0000% 0.0000% 0.0000% 0.0000%
SaltPepper 0.0048% 0.0033% 0.0428% 0.0226% 0.0132%
GaussianNoise 0.0032% 0.0208% 0.5491% 0.0108% 0.0211%
Average 0.0039% 0.0035% 0.0777% 0.0045% 0.0047%
SimSwap 0.2175% 0.2139% 0.9971% 0.0713% 0.0287%
GANimation 0.0015% 0.0000% 0.0038% 0.0013% 0.0000%
StarGAN (Male) 0.0247% 0.0067% 0.1207% 0.0069% 0.0051%
StarGAN (Young) 0.0138% 0.0032% 0.1069% 0.0044% 0.0019%
StarGAN (BlackHair) 0.0103% 0.0017% 0.1128% 0.0031% 0.0022%
StarGAN (BlondHair) 0.0119% 0.0027% 0.1249% 0.0063% 0.0013%
StarGAN (BrownHair) 0.0076% 0.0089% 0.1416% 0.0097% 0.0064%
Average 0.0410% 0.0339% 0.2297% 0.0147% 0.0065%
Table IV
ABLATION RESULTS FOR PSNR, SSIM, AND DETECTOR’S BIT ERROR RATE (%), WHERE ‘×’ DENOTES NON-USAGE AND ‘$\sqrt{}$’ DENOTES USAGE.
Case1 Case2 Case3 Case4 Proposed
Joint Sub-bands $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Attention module $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
frequency domain $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
SC-GNN $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
PSNR $\uparrow$ 45.2321 43.7024 49.0436 42.8801 47.5555
SSIM $\uparrow$ 0.9956 0.9823 0.9955 0.9517 0.9957
Identity 0.0000% 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.0000% 0.0000% 0.0152% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0224% 0.0000% 0.0000%
GaussianBlur 0.0017% 0.0027% 0.0064% 0.0013% 0.0021%
MedianBlur 0.0011% 0.0024% 0.0019% 0.0009% 0.0017%
Dropout 0.0000% 0.0000% 0.0000% 0.0000% 0.0000%
SaltPepper 0.0078% 0.0178% 0.0123% 0.0083% 0.0156%
GaussianNoise 0.0149% 0.0128% 0.7410% 0.0132% 0.0327%
Average 0.0032% 0.0045% 0.0999% 0.0030% 0.0065%
SimSwap 45.4832% 46.3146% 46.7812% 46.1530% 46.7845%
GANimation 45.4851% 46.3144% 46.7856% 46.1534% 46.7849%
StarGAN (Male) 45.4835% 46.3132% 46.7837% 46.1227% 46.7828%
StarGAN (Young) 45.4842% 46.3187% 46.7836% 46.1510% 46.7828%
StarGAN (BlackHair) 45.4811% 46.3131% 46.7842% 46.1574% 46.7845%
StarGAN (BlondHair) 45.4896% 46.3145% 46.7813% 46.1533% 46.7811%
StarGAN (BrownHair) 45.4842% 46.3149% 46.7811% 46.2061% 46.7828%
Average 45.4844% 46.3148% 46.7830% 46.1567% 46.7833%
IV-D Ablation Study
In order to evaluate the contribution of Joint Sub-bands, Attention Module, Frequency Domain Embedding, and GNN in this method, we conducted ablation experiments on both Tracer and Detector. The experiments were conducted on images with a resolution of 128×128. (a) Adjust the sub-bands usage of Tracer and Detector: In Tracer, watermark embedding was reduced from 1,3,4,6 sub-bands to 1,3, while in Detector, the detection was expanded from 1,3 sub-bands to 1,3,4,6. This allows us to evaluate the influence of joint sub-band information redundancy on watermark recovery and deepfake detection. (b) Remove the attention mechanism in the process of watermark embedding to analyze its influence on visual quality; (c) When the DT-CWT frequency domain embedding is removed, the watermark is directly embedded in the U channel. (d) Removing SC-GNN to evaluate its impact on the visual quality of the watermarked image. Tables 3 and 4 present the results of the ablation study.
It can be found that: (a) From Case 1, the joint sub-band significantly affects the robustness and detection ability of the watermark. When Tracer only extracts the watermark in 1 and 3 sub-bands, the BER increases obviously, indicating that reducing sub-bands weakens the stability of the watermark. However, when the Detector is extended to 1, 3, 4 and 6 sub-bands, the BER decreases, and the extra sub-bands enhance the redundancy of the watermark, making its detection more stable under common distortion and deepfake forgery. (b) From Case 2, it can be seen that the PSNR and SSIM of the watermark image decrease after the attention mechanism is removed, which indicates that the attention mechanism has optimized the watermark embedding. Although BER has not changed much, the invisibility and visual quality of the watermark have obviously decreased, which verifies the importance of attention mechanism in optimizing visual quality. (c) From Case 3, it can be seen that the robustness of the watermark is significantly reduced after the frequency domain embedding is removed. Although PSNR and SSIM remain at a high level, BER increases significantly under common distortion and deepfake forgery. In addition,the perfect reconstruction, directional selectivity and approximate translation invariance of DT-CWT make frequency domain embedding more resistant to distortion and forgery, which verifies its key role in improving watermark robustness. (d) From Case 4, it can be seen that after removing SC-GNN, the PSNR and SSIM of the watermark image decrease, although the BER changes little, indicating that SC-GNN plays an important role in optimizing the visual quality of the watermark image. SC-GNN improves the invisibility of the watermark and reduces perceptual artifacts by extracting the structural features of the image, which verifies the key role of SC-GNN in optimizing the visual quality of the watermark.
V CONCLUSION
In this paper, a watermarking framework based on frequency domain embedding and Structural Consistency Graph Neural Network (SC-GNN) is proposed. The robustness is improved by Dual-Tree Complex Wavelet Transform (DT-CWT) transform, and the detection ability of deepfake is enhanced by combining sub-band information. In addition, the framework leverages dense connectivity to enhance feature reuse and robustness. SC-GNN further optimizes the visual quality of watermark embedding, reduces perceptual artifacts, and improves the invisibility of watermark by attention mechanism. Experimental results show that this method can stably extract watermark and accurately detect tampering in various distorted and forged scenes. Future work will further improve its generalization ability and optimize its adaptability to complex distortion and unknown deepfake technology.